OpenStudy (anonymous):

A number is five times another number. When the numbers are added, their sum is 24. What are the two numbers?

OpenStudy (radar):

Let x = one number Let 5x = other number 5x+x = 24 6x=24 x=4 5x=20 The numbers are 4 and 20

OpenStudy (anonymous):

5time 6 is 30 how does it work to equal 24?

OpenStudy (radar):

Did you follow that antonio0111?

OpenStudy (radar):

5 times 6 is 30, no one is saying it is 24.

OpenStudy (radar):

However 6 times 4 is 24.

OpenStudy (anonymous):

yeah with help I got it thank it

OpenStudy (anonymous):

the problem seems to be that radar has got only one number(variable) in the equation and in the end there are 2 solutions

OpenStudy (radar):

Good, keep practicing, and good luck with your studies.

OpenStudy (anonymous):

I got one more for you if you don't mind, I get what you mean better then most Radar

OpenStudy (radar):

It is true I could of used two variables and maybe it would of been clearer, but I like less variables.

OpenStudy (radar):

We are here to help antonio0111.

OpenStudy (anonymous):

11 - 8 ÷ (1/3 + 1/5)

OpenStudy (anonymous):

That is where I say 3/15 +5/15=8/15

OpenStudy (anonymous):

Am I right?

OpenStudy (radar):

Clear the parenthesis: (5/15+3/15)= 8/15 Now the problem looks like this \[11-8\div8/15\] Remember the order of operations......division has priority over addition so now: 11 -(8 x15/8) 11 - 15=-4

OpenStudy (radar):

Yes you added the part in the parenthesis correctly.